# Reformulating Domain Adaptation of Large Language Models as Adapt-Retrieve-Revise

### v1

@@ Abstract
While large language models (LLMs) like GPT-4 have recently demonstrated astonishing zero-shot capabilities in general domain tasks, they often generate content with hallucinations in specific domains such as Chinese law, hindering their application in these areas. This is typically due to the absence of training data that encompasses such a specific domain, preventing GPT-4 from acquiring in-domain knowledge. A pressing challenge is that it’s not plausible to continue training LLMs of such scale on in-domain data.
@@
This paper introduces a simple and effective domain adaptation framework for GPT-4 by reformulating generation as an adapt-retrieve-revise process. The initial step is to adapt an affordable 7B LLM to the target domain by continuing learning on in-domain data. When solving a task, we leverage the adapted LLM to generate a draft answer given a task query. Then, the draft answer will be used to retrieve supporting evidence candidates from an external in-domain knowledge base. Finally, the draft answer and retrieved evidence are concatenated into a whole prompt to let GPT-4 assess the evidence and revise the draft answer to generate the final answer.
@@
Our proposal combines the advantages of the efficiency of adapting a smaller 7B model with the evidence-assessing capability of GPT-4 and effectively prevents GPT-4 from generating hallucinatory content. In the zero-shot setting of four Chinese legal tasks, our method improves accuracy by 33.3% compared to the direct generation by GPT-4. When compared to two stronger retrieval-based baselines, our method outperforms them by 15.4% and 23.9%. Our code will be released  [^1] The link will be updated later. The domain adapted 7B LLM will be also released..
1 Introductions
@@
Recent large language models (e.g., GPT-4) bring remarkable improvements in various general domain NLP tasks (Brown et al., 2020a; OpenAI, 2023; Thoppilan et al., 2022; Chowdhery et al., 2022; Rae et al., 2022; Hoffmann et al., 2022). However, in specific domains such as the Chinese legal domain, the performance of such general LLMs still lags considerably behind. We show a real example of Chinese LegalQA (Chen et al., 2023) on the left of Figure 1, which requires the model to generate the corresponding legal provision (i.e., the law name and the clause index) and the rationale for the judgment, given a brief case description as the query.
@@
We initialize the research with a preliminary examination of utilizing GPT-4 to address the Chinese LegalQA task, which involves responding with a law clause relevant to a given query case. Figure 1 reveals the extremely low recall of directly prompting the query case to ask GPT-4 to generate the corresponding law clause. Though the generated answers are grammatically fluent, they often consist of non-logical content, factual mistakes, and fail to refer to the correct legal provision (also known as ”hallucination”). For example, in Figure 2, the direct generation of GPT-4 seems to be logical but actually has no clues in the Chinese laws. A potential reason is the insufficient Chinese legal domain text during pretraining, leading to a lack of domain knowledge acquisition, and consequently generating hallucinatory content.
@@
For the LLMs with the scale of GPT-4, it’s generally not feasible for researchers to do learning-based adaptation. The enormous model size could make the cost of continual learning extremely high, and meanwhile, the access functions are often limited by APIs. Therefore, recent work (Lewis et al., 2020b) introduces retrieval-based methods that first use the given query to retrieve relevant evidence candidates from the external domain-specific knowledge base or the internet and then concatenate the query and the evidence candidates into the prompt. GPT-4 could implicitly validate the relevance between the query and the evidence, as well as the correctness of the evidence, before producing a generation. In our replicated results, the retrieval-based method improves the LegalQA recall from 14.4% (direct generation) to 36.0%. This suggests that even though GPT-4 may not generate domain content, it possesses sufficient evidence-assessing capacity to select the correct evidence from candidates. Nevertheless, the retrieval module is limited by the capability of representation mapping from query to evidence and is also influenced by the domain issue, leading to a decline in search quality. GPT-4 still produces hallucinations in responses as the middle answer in Figure 2.
@@
On the other hand, with the rapid development of open LLMs led by LLaMA (Touvron et al., 2023), it becomes affordable to continually train an open LLM tailored to your demands on sufficient in-domain texts, resulting in a domain-adapted LLM. We therefore conduct the second examing of continually training Baichuan 7B (Baichuan-inc, 2023), a Chinese foundation model, on over 50B Chinese legal data. Its performance significantly outperforms GPT-4 and even surpasses the retrieval-based GPT-4 generation on Chinese LegalQA. Hallucinations caused by the lack of domain knowledge are largely reduced but not completely solved. As shown in the right answer in Figure 2, adapted LLM generates generally correct responses but still makes errors in certain words. Although the law name is correct and the rationale part is reasonable, the clause index is a hallucinatory generation which raises the difficulty in anchoring target clauses. We argue that these fails are accordingly caused by the limited capability of a 7B size to understand the query and the evidence.
@@
Figure 1: Left: A real example of Chinese LegalQA. The square brackets and subscripts are offered for the purpose of clear demonstration, not actually exist in the ground-truth answer or generation. Right: The recall on the LegalQA dataset.
The rest sections of the paper anchor the Chinese legal domain and comprehensively validate the effectiveness of our proposal. In Section 2, we explain each stage of our adapt-retrieve-revise method and elaborate on the implementation details. In Section 3 and 4, we conduct the experiments and the result analysis on four Chinese legal domain tasks. The experimental results show that our proposal significantly improves the direct generation and the retrieval-based generation baselines. In the final Section 6, we elicit the conclusion and future work. To the best of our knowledge, this is the first study to examine the zero-shot performance of LLMs on four Chinese legal benchmarks.
Figure 2: Examples of hallucinations of various models. Red denotes the content containing hallucinations. The ground-truth answer is shown in Figure 1
@@ 2 Methodology
Figure 3: Overview of our proposed method.
2.1 Generate the draft answer by the domain-adapted LLM
Recently, various Chinese pre-trained LLMs have occurred, such as Baichuan 7B, Qwen 7B (Alibaba-inc, 2023), etc. In this paper, we chose Baichuan-7B as the base model for our proposal. For the Chinese legal raw corpora, we build the continual learning data from two open sources:
@@ 2.1 Generate the draft answer by the domain-adapted LLM
  • Chinese judgments online (https://wenshu.court.gov.cn/) is the largest online publication platform for legal documents in China. The platform contains judicial documents from courts at all levels, covering various legal fields such as civil, criminal, administrative, and enforcement. We believe these documents contain knowledge for LLMs to understand the usage of laws in various scenarios, and we collect 100M documents from this platform as the training data.
In total, we trained 50B tokens with the input length limit of 16,384 and the batch size of 256 on 32 A100 GPUs, and the time-consuming is 167 hours. After continual learning, we subsequently supervised fine-tuning our model on 70K instruction examples, including 52K GPT-4 self-instruct Chinese data (Peng et al., 2023) and 18K legal instructions (See Appendix A.1) for the alignment.
Then, given an input query, we will first prompt the trained 7B legal LLM to generate the draft answer, which will be used in the next step. For the prompt, we add the instruction ”Please provide evidence in the Chinese law” at the end of the query to enforce the model to generate related law clauses, as in Figure 3.
@@ 2.2 Answer-based Evidence Retrieval
For the $i$-th paragraph $p_{i}$, we construct the key-value pair $(\displaystyle{\bm{p}}_{i},p_{i})$ where the key ${\bm{p}}_{i}$ is the representation obtained from the sentence embedding model and the value $p_{i}$ denotes the paragraph. In our experiment, we utilize Multilingual-E5-large (Wang et al., 2022), a Roberta-based (Liu et al., 2019) sentence embedding model that achieves robust performance on various tasks. The memory $(\mathcal{K},\mathcal{V})=\{(\displaystyle{\bm{p}}_{i},p_{i})|p_{i}\in\mathcal{KB}\}$ is thus the set of all key-value pairs constructed from all the paragraphs in the external knowledge base $\mathcal{KB}$.
@@
Given the generated draft answer $d$, the sentence embedding model E5 outputs its representation ${\bm{h}}_{Answer}$. We then query the constructed knowledge bank with ${\bm{h}}_{Answer}$ to retrieve its $k$ nearest neighbors $E$ according to a distance function by $L^{2}$ distance.
@@ 2.3 GPT-4 Revision
3 Experiments
We conducted a series of experiments to compare our adapt-retrieve-revise method to the baselines of direct generation and retrieval-based generation on various Chinese legal benchmarks. We find that our proposal robustly improves performance on all the tasks. We show the task settings and experimental results in this section.
3.1 Chinese Legal Domain Tasks
We evaluated our Adapt-Retrieve-Revise method on a diversity of tasks in the zero-shot setting, and we divided these tasks into three categories by the legal knowledge base for retrieval:
@@ 3.1 Chinese Legal Domain Tasks
  • Similar Case Retrieval (Ma et al., 2021) is the task that, given a query legal scenario as the input, we aim at selecting similar Legal Judgement Documents from the $100$ candidates. We conducted this experiment to assess the reliability of our proposed retrieval method in Section 4.1.2.
Due to the cost of GPT-4 API and the human evaluation, we randomly sampled a subset of $250$ test examples for each task of LCR, CP, LegalQA, and JEC-QA.
@@ 3.2 Evaluation Metrics
Since generative models produce diverse formats in the output and the Chinese legal domain has its own features, checking the evaluation metrics in the experiments is crucial. For tasks LCR, CP, and LegalQA, our metric is the recall of whether the title of the ground-truth law clause is included in the generated answer. This is because, in real-world applications, with the correct title, the contents of the law clause can be easily revised by the rule-based system, indicating that the title is more important than the content.
For the JEC-QA task, we use accuracy as the metric, but controlling the output into an identical format for automatic evaluation is difficult, especially for the 7B LLM that has not been fine-tuned on the JEC-QA task. We select human evaluation to ensure the accuracy of our evaluation.
@@
3.3 Main Results
Table 1: Main Results on four Chinese legal datasets.
Models Retrieval Revisor Chinese law articles Textbooks Avg.
LCR CP LegalQA JEC-QA
Direct Generation
GPT-4 - - 67.6 71.2 14.4 36.2 47.4
7B legal LLM - - 88.4 84.0 48.8 39.8 65.3
Retrieval-based Generation
GPT-4 Query-based - 74.4 75.2 36.0 41.6 56.8
7B legal LLM Query-based - 87.6 82.4 50.2 40.8 65.3
Adapt-Retrieve-Revise
7B legal LLM Answer-based 7B legal LLM 88.4 83.6 49.0 40.2 65.3
7B legal LLM (ours) Answer-based GPT-4 96.4 87.8 72.4 66.2 80.7
We provide the main results as in Table 1. Generally, we compare our proposed method with direct generations and retrieval-based generations using the query, showing that our method outperforms all baselines by a large margin. Our main results also provide some ablation results.
First, our 7B legal LLM significantly beats GPT-4, and even compared with the retrieval-based generation of GPT-4, the 7B legal LLM still outperforms on three tasks and has competitive results on the JEC-QA task, indicating that our continual learning on Chinese legal raw corpora shows a fast and effective domain adaptation on various legal tasks.
Then, considering the results of GPT-4 and the GPT-4 retrieval-based generation, we find that after providing evidence of related legal knowledge, GPT-4 can improve its responses significantly ($+9.92\%$). This indicates that the retrieval-based method is a proper way to reduce hallucinations caused by the lack of domain knowledge, and owing to the robust evidence-assessing capacity, GPT-4 can adapt to the Chinese legal domain well with convincing evidence available.
In our final experiment, using the draft answers generated by the 7B legal model for retrieval and revision, the performance significantly exceeded two query-based retrieval baselines by large margins of 15.4% and 23.9%. It’s worth noting that the improvement here comes from both the enhanced answer-based retrieval quality and the revision setup. In the subsequent ablation study, we will further examine the quantified improvement of the retrieval quality through an additional retrieving task.
4 Ablation Studies & Further Analysis
4.1 Analysis of Retrieval Methods

### v3

@@ Abstract
While large language models (LLMs) like GPT-4 have recently demonstrated astonishing zero-shot capabilities in general domain tasks, they often generate content with hallucinations in specific domains such as Chinese law, hindering their application in these areas. This is typically due to the absence of training data that encompasses such a specific domain, preventing GPT-4 from acquiring in-domain knowledge. A pressing challenge is that it’s not plausible to continue training LLMs of the GPT-4’s scale on in-domain data.
@@
This paper introduces a simple yet effective domain adaptation framework for GPT-4 by reformulating generation as an adapt-retrieve-revise process. The initial step is to adapt an affordable 7B LLM to the Chinese legal domain by continuing learning in-domain data. When solving an in-domain task, we leverage the adapted LLM to generate a draft answer given a task query. Then, the draft answer will be used to retrieve supporting evidence candidates from an external in-domain knowledge base. Finally, the draft answer and retrieved evidence are concatenated into a whole prompt to let GPT-4 assess the evidence and revise the draft answer to generate the final answer.
@@
Our proposal combines the advantages of the efficiency of adapting a smaller 7B model with the evidence-assessing capability of GPT-4 and effectively prevents GPT-4 from generating hallucinatory content. In the zero-shot setting of four Chinese legal tasks, our method improves the average score by $+33.6$ points, compared to GPT-4 direct generation. When compared to two stronger retrieval-based baselines, our method outperforms them by $+17.0$ and $+23.5$. The code for training our model is here: 111 https://github.com/YukinoWan/Adapt-Retrive-Revise.
Reformulating Domain Adaptation of Large Language Models as Adapt-Retrieve-Revise: A Case Study on Chinese Legal Domain
Zhen Wan Kyoto University zhenwan.nlp@gmail.com                        Yating Zhang Alibaba Health yatingz89@gmail.com                        Yexiang Wang Alibaba Quark yexiang_w@163.com
Fei Cheng Kyoto Univeristy feicheng@i.kyoto-u.ac.jp                        Sadao Kurohashi Kyoto University kuro@i.kyoto-u.ac.jp
1 Introduction
Figure 1: Left: A real translated example of Chinese LegalQA. The square brackets and subscripts are offered for the purpose of clear demonstration, not actually exist in the ground-truth answer or generation. Right: Models’ F1 scores on the LegalQA dataset.
@@ 1 Introduction
Recent large language models like GPT-4 bring remarkable improvements in various general domain NLP tasks (Brown et al., 2020a; OpenAI, 2023a; Thoppilan et al., 2022; Chowdhery et al., 2022; Rae et al., 2022; Hoffmann et al., 2022). However, in specific domains such as the Chinese legal domain, the performance of such general LLMs still lags considerably behind. We show a real example of Chinese LegalQA (Chen et al., 2023) on the left of Figure 1, which requires the model to generate the corresponding legal provision (i.e., the law name and the clause index) and the rationale for the judgment, given a brief case description as the query.
@@
We initialize the research with a preliminary examination of utilizing GPT-4 to address the Chinese LegalQA task, which involves responding with a law clause relevant to a given query case. Figure 1 reveals the extremely low performance (F1 $12.0$) of directly prompting the query case to ask GPT-4 to generate the corresponding law clause. Though the generated answers are grammatically fluent, they often consist of non-logical content, factual mistakes, and fail to refer to the correct legal provision (also known as "hallucination"). For example, in the left answer of Figure 2, the direct generation of GPT-4 seems to retell the case description but fails to point out the corresponding clause. A potential reason is the insufficient Chinese legal domain text during pretraining, leading to a lack of domain knowledge acquisition, and consequently generating hallucinatory content.
@@
For the LLMs with the scale of GPT-4, it’s generally not feasible for researchers to do learning-based adaptation. The enormous model size could make the cost of continual learning extremely high, and meanwhile, the access functions are often limited by APIs. Therefore, recent work (Lewis et al., 2020b; Yu et al., 2023; Shuster et al., 2021; Ma et al., 2023) introduces retrieval-based methods that first use the given query to retrieve relevant evidence candidates from the external domain-specific knowledge base or the internet and then concatenate the query and the evidence candidates into the prompt. GPT-4 could implicitly validate the relevance between the query and the evidence, as well as the correctness of the evidence, before producing a generation.
Our replicated retrieval-based method improves the LegalQA F1 from the $12.0$ points of direct generation to $33.1$ as shown in Figure 1. It indicates that even though GPT-4 may not generate domain content, it possesses sufficient evidence-assessing capacity to select the correct evidence from candidates. Nevertheless, the retrieval module is limited by the capability of representation mapping from query to evidence and is also influenced by the domain issue, leading to a decline in search quality. GPT-4 still produces hallucinations in responses as the middle answer in Figure 2.
@@
On the other hand, with the rapid development of open LLMs led by LLaMA (Touvron et al., 2023), it becomes affordable to continually train an open LLM tailored to your demands on sufficient in-domain texts, resulting in a domain-adapted LLM. We therefore conduct the second examing of continually training Baichuan 7B (Baichuan-inc, 2023), a Chinese foundation model, on over 50B token Chinese legal data. Its performance (the green bar in Figure 1) even surpasses the retrieval-based GPT-4 generation on Chinese LegalQA. Hallucinations caused by the lack of domain knowledge are largely reduced but not completely solved. As shown in the right answer in Figure 2, adapted LLM generates generally correct responses but still makes errors in certain words. Although the law name is correct and the rationale part is reasonable, the clause index is a hallucinatory generation which raises the difficulty in anchoring target clauses. We argue that these fails are accordingly caused by the limited capability of a 7B size to memorize the knowledge accurately.
@@
The rest sections of the paper anchor the Chinese legal domain and comprehensively validate the effectiveness of our proposal. In Section 2, we explain each component of our adapt-retrieve-revise method and elaborate on the implementation details. In Section 3 and 5, we conduct the experiments and the result analysis on four Chinese legal domain tasks. The experimental results show substantial improvements against the direct generation and the retrieval-based generation baselines. In the final Section 7, we elicit the conclusion and future work. To the best of our knowledge, this is the first study to examine the zero-shot performance of LLMs on four Chinese legal benchmarks.
Figure 2: Examples of hallucinations of various models. Red denotes the content containing hallucinations. The ground-truth answer refers to the left case in Figure 1.
Figure 3: Overview of our proposed method. The example and prompt are translated from Chinese to English for the demonstration purpose.
@@ 2 Methodology
2.1 Draft Answer Generation by the Domain-adapted LLM
This step could be actually flexible to save the effort of training a LLM yourself. For instance, an off-the-shelf Chinese legal domain LLM can be used for generating draft answer. However, most of such LLMs have been fine-tuned on various in-domain evaluation tasks. The potential data leakage could lead to unreliable evaluation, especially in a zero-shot setting. Therefore, in this paper, we adapt an open 7B LLM to the Chinese legal domain by continual learning on the domain data. We collect the training data from the following two open Chinese legal sources:
@@ 2.1 Draft Answer Generation by the Domain-adapted LLM
  • Chinese Judgments Online (https://wenshu.court.gov.cn/) is the largest online publication platform for legal documents in China. The platform contains judicial documents from courts at all levels, covering various legal fields such as civil, criminal, administrative, and enforcement. Such documents contain knowledge for LLMs to understand the usage of laws in various scenarios.
During the inference, given an input query, we will first prompt the trained 7B legal LLM to generate the draft answer, which will be used in the next step. For the prompt, we add the instruction ”Please provide evidence in the Chinese law" at the end of the query to enforce the model to generate related law clauses, as in Figure 3.
@@ 2.2 Answer-based Evidence Retrieval
For the $i$-th paragraph $p_{i}$, we construct the key-value pair $(\mathbf{p_{i}},p_{i})$ where the key $\mathbf{p_{i}}$ is the representation obtained from the sentence embedding model and the value $p_{i}$ denotes the paragraph. The memory $(\mathcal{K},\mathcal{V})=\{(\mathbf{p_{i}},p_{i})|p_{i}\in\mathcal{KB}\}$ is thus the set of all key-value pairs constructed from all the paragraphs in the external knowledge base $\mathcal{KB}$.
@@
Given the generated draft answer $d$, the sentence embedding model outputs its representation $\mathbf{h_{Answer}}$. We then query the constructed knowledge bank with $\mathbf{h_{Answer}}$ to retrieve its $k$ nearest neighbors $E$ according to a distance function by $L^{2}$ distance.
@@ 2.3 GPT-4 Revision
3 Experiment Settings
We conducted a series of experiments to compare our adapt-retrieve-revise method to the baselines of direct generation and retrieval-based generation on various Chinese legal benchmarks. We show the model details and the task settings in this section.
3.1 Model Settings
Details of training 7B legal LLM:
[…]
Retriever setting:
We utilize Multilingual-E5-large (Wang et al., 2022), a Roberta-based (Liu et al., 2019) sentence embedding model that achieves robust performance on various tasks. We also compare with other retrieval modules in Appendix A.2.
GPT-4 setting:
[…]
3.2 Evaluation of Chinese Legal Tasks
We evaluated our Adapt-Retrieve-Revise method on a diversity of tasks with different knowledge base for retrieval in the zero-shot setting:
@@ 3.2 Evaluation of Chinese Legal Tasks
  • Similar Case Retrieval (Ma et al., 2021) is the task, given a query legal scenario as the input, we aim at selecting similar Legal Judgement Documents from the $100$ candidates. We conducted this experiment to assess the reliability of our proposed retrieval method in Section 5.1.2.
Generator Retriever Revisor Chinese law clauses Textbooks Avg.
LCR CP LegalQA JEC-QA
F1 (Rec.) F1 (Rec.) F1 (Rec.) Acc.
Direct Generation
GPT-4 - - 61.7 (67.6) 70.7 (71.2) 12.0 (14.4) 36.2 45.1
LawGPT-7B Song (2021) - - 19.3 (25.3) 33.3 (34.6) 10.3 (16.6) 27.4 22.6
ChatLaw-13B Cui et al. (2023a) - - 25.6 (27.6) 43.6 (49.6) 14.6 (18.3) 31.8 28.9
7B legal LLM - - 83.0 (88.4) 82.9 (84.0) 41.9 (48.8) 39.8 61.9
Retrieval-based Generation
GPT-4 Query-based - 72.0 (74.4) 74.0 (75.2) 33.1 (36.0) 41.6 55.2
7B legal LLM Query-based - 77.3 (87.6) 81.3 (82.4) 47.4 (50.2) 40.8 61.7
Adapt-Retrieve-Revise
7B legal LLM Answer-based 7B legal LLM 84.1 (88.4) 82.7 (83.6) 47.3 (49.0) 40.2 63.6
7B legal LLM (ours) Answer-based GPT-4 90.6 (96.4) 86.9 (87.8) 71.1 (72.4) 66.2 78.7
Table 1: Zero-shot Results on four Chinese legal datasets. “Rec.” denotes recall, “Acc.” denotes accuracy, and “Avg.” is computed by the F1 and accuracy scores of all four tasks
Due to the cost of GPT-4 API and the human evaluation, we randomly sampled a subset of $250$ test examples for each task of LCR, CP, LegalQA, and JEC-QA. Please refer to Table 2 for the statistics.
Datasets CP LCR LegalQA JEC-QA
# test 965, 219 965, 219 1,000 13,341
Table 2: Statistics of datasets.
@@ 3.3 Evaluation Metrics
Since generative models produce diverse formats in the output and the Chinese legal domain has its own features, checking the evaluation metrics in the experiments is crucial. For tasks LCR, CP, and LegalQA, our metric is the Micro F1 and the Recall of whether the title of the ground-truth law clause is included in the generated answer. This is because, in real-world applications, with the correct title, the contents of the law clause can be easily revised by the rule-based system, indicating that the title is more important than the content.
For the JEC-QA task, we use accuracy as the metric, but controlling the output into an identical format for automatic evaluation is difficult, especially for the 7B LLM that has not been fine-tuned on the JEC-QA task. We select human evaluation to ensure its accuracy by manually comparing the model output and the gold answer provided by the dataset.
@@
4 Main Experimental Results
We provide the main results as in Table 1. As we claimed before, this is the first work to targeting the zero-shot LLM performance on Chinese legal domain tasks. Generally, we compare our adapt-retrieve-revise proposal with baselines of direct generations and retrieval-based generations using the query, showing that our method outperforms all baselines by substantial margins. Our main results also provide some ablation results.
We first observe the effectiveness of domain adaption. Our 7B legal LLM significantly beats GPT-4, and even compared with the retrieval-based generation of GPT-4, the 7B legal LLM still outperforms on three tasks and has competitive results on the JEC-QA task, indicating that our continual learning on Chinese legal raw corpora shows a fast and effective domain adaptation on various legal tasks. However, related work (LawGPT and ChatLaw) fails to benefit largely from the continual training due to the much less data used in training, and their base models are multilingual Llama.
Then, considering the results of GPT-4 and the GPT-4 retrieval-based generation, we find that after providing evidence of related legal knowledge, GPT-4 can improve its responses significantly ($+10.1$ points). This indicates that the retrieval-based method is a proper way to reduce hallucinations caused by the lack of domain knowledge, and owing to the robust evidence-assessing capacity, GPT-4 can adapt to the Chinese legal domain well with convincing evidence available.
@@ 4 Main Experimental Results
In our final experiment, using the draft answers generated by the 7B legal model for retrieval and revision, the performance significantly exceeded two query-based retrieval baselines by large margins of $+17.0$ and $+23.5$ points. It’s worth noting that the improvement here comes from both the enhanced answer-based retrieval quality and the revision setup. In the subsequent ablation study, we will further examine the quantified improvement of the retrieval quality through an additional retrieving task.
5 Further Analysis
5.1 Analysis of Retrieval Methods
The previous section has demonstrated the step-wise effectiveness of our adapt-retrieve-revise proposal. Nevertheless, the retriever, as a key component that directly affects the quality of evidence, how its variations would impact the final performance is a crucial research question to investigate.